OpenStudy (anonymous):

Which of the following is the slope between the points (-1, -5) and (7, -3)? 4 1/4 -1/4 -4

OpenStudy (campbell_st):

m = differene in y/ difference in x m = (-3 - -5)/7 - - 1) m = 2/8 m = 1/4